OpenStudy (anonymous):

which is the best estimate for 1/2 % of 385

OpenStudy (ash2326):

We have \[\frac{1}{2} \%\ of\ 385\] It's equivalent to \[\frac{1}{2} \%\ \times 385\] \[\frac{1}{2 \times 100} \times 385\] \[\frac{1}{\cancel {2} \times 100} \times \cancel{385} 192.5\] we get \[1.925\]

OpenStudy (anonymous):

0.005*385

OpenStudy (anonymous):

If you were to *estimate* it, the best way would be to first round 385 off to 400 (we are estimating here after all), and then take 1% of that, which is 4. Then you half 4 to get 2.

OpenStudy (amistre64):

385/100 = 1% 385/100 -------- = 1/2 % 2

OpenStudy (anonymous):

\[385\approx400\] 400*1%=4 4/2=2